# Monogamy of nonlocality from multipartite information causality

### v1

@@ Abstract
$p_{\tilde{\mathbb{P}}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}}(a,b,e|x,y,z)=\frac{1}{2}p_{\mathbb{P}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}}(a,b,e|x,y,z)+\frac{1}{2}p_{\bar{\mathbb{P}}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}}(a,b,e|x,y,z),$ (37)
@@
$p_{\mathbb{P}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}}(a,b,e|x,y,z)$ $=\alpha p_{\text{PR}_{\mathcal{A}\mathcal{B}}\otimes\text{WN}_{\mathcal{E}}}(a,b,e|x,y,z)+\gamma p_{\text{PR}_{\mathcal{B}\mathcal{E}}\otimes\text{WN}_{\mathcal{A}}}(a,b,e|x,y,z)$ (38)
$\ \ \ \ \ \ \ +(1-\alpha-\gamma)p_{\text{WN}_{\mathcal{A}}}(a|x)p_{\text{WN}_{\mathcal{B}}}(b|y)p_{\text{WN}_{\mathcal{E}}}(e|z),$ (39)
@@
$=\alpha\frac{1}{4}\delta_{a\oplus b,xy}+\gamma\frac{1}{4}\delta_{b\oplus e,yz}+(1-\alpha-\gamma)1/8,$ (40)
where for the first equality we have used the facts, $\frac{1}{2}(L_{\mathcal{B}}+\bar{L}_{\mathcal{B}})=\text{WN}_{\mathcal{B}}$, $\frac{1}{2}(L_{\mathcal{E}}+\bar{L}_{\mathcal{E}})=\text{WN}_{\mathcal{E}}$, $\frac{1}{2}(\mathbb{P}^{(\alpha,\gamma)}_{\mathcal{A}}+\bar{\mathbb{P}}^{(\alpha,\gamma)}_{\mathcal{A}})=\text{WN}_{\mathcal{A}}$, $\frac{1}{2}(\mathbb{P}^{(\alpha,\gamma)}_{\mathcal{B}}+\bar{\mathbb{P}}^{(\alpha,\gamma)}_{\mathcal{B}})=\text{WN}_{\mathcal{B}}$, and $\frac{1}{2}(\mathbb{P}^{(\alpha,\gamma)}_{\mathcal{E}}+\bar{\mathbb{P}}^{(\alpha,\gamma)}_{\mathcal{E}})=\text{WN}_{\mathcal{E}}$.
Notice that the values of the CHSH expression, $\beta(\mathcal{A},\mathcal{B})$ and $\beta(\mathcal{B},\mathcal{E})$ still remain unaltered as they only depend on the coefficients $(\alpha,\gamma)$ (27) (28). Now, to complete the proof, all that remains is to demonstrate that if $\mathbb{P}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}$ satisfies the bipartite (7) and/or multipartite (8) $\mathcal{IC}$ criteria for some $(\alpha,\gamma)$, then the corresponding distribution ~$\tilde{\mathbb{P}}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}$ for the same $(\alpha,\gamma)$ also satisfies the bipartite (7) and/or multipartite (8) formulations of $\mathcal{IC}$. To this end, we invoke convexity of mutual information on probability distributions, which yields, for the bipartite $\mathcal{IC}$ criterion (7),
$\sum_{i=1}^{n}I_{{\tilde{\mathbb{P}}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}}}(X_{i}:G_{i},M^{\prime})+\sum_{i=2}^{n}I_{\tilde{\mathbb{P}}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}}(X_{1}:X_{i}|G_{i},M^{\prime})$ $\leq\frac{1}{2}\left(\sum_{i=1}^{n}I_{{{\mathbb{P}}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}}}(X_{i}:G_{i},M^{\prime})+\sum_{i=2}^{n}I_{{\mathbb{P}}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}}(X_{1}:X_{i}|G_{i},M^{\prime})\right)$
$\ \ \ \ \ \ \ +\frac{1}{2}\left(\sum_{i=1}^{n}I_{{\bar{\mathbb{P}}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}}}(X_{i}:G_{i},M^{\prime})+\sum_{i=2}^{n}I_{\bar{\mathbb{P}}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}}(X_{1}:X_{i}|G_{i},M^{\prime})\right),$
$\leq\sum_{i=1}^{n}I_{{{\mathbb{P}}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}}}(X_{i}:G_{i},M^{\prime})+\sum_{i=2}^{n}I_{{\mathbb{P}}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}}(X_{1}:X_{i}|G_{i},M^{\prime}),$ (41)
where for the first inequality we have invoked the fact that the mutual information is a convex function of the probability distributions over the protocol-specific variables $\{X_{i}\},\{G_{i}\},M^{\prime}$ which in turn are convex linear functions of the underlying tripartite no-signaling distributions ${\mathbb{P}}^{(\alpha,\gamma)},\bar{\mathbb{P}}^{(\alpha,\gamma)},\tilde{\mathbb{P}}^{(\alpha,\gamma)}$, while the second inequality follows from (35), and similarly for the multipartite $\mathcal{IC}$ criterion (8),
$\sum_{k,i}I_{{\tilde{\mathbb{P}}_{\mathcal{A}\mathcal{B}\mathcal{E}}}}(X_{i}^{k}:X_{i}^{1},\dots,X_{i}^{k-1},X_{i}^{k+1},\dots,X_{i}^{N-1},\mathbf{M^{\prime}},G_{i})$ $\leq\frac{1}{2}\sum_{k,i}I_{{{\mathbb{P}}_{\mathcal{A}\mathcal{B}\mathcal{E}}}}(X_{i}^{k}:X_{i}^{1},\dots,X_{i}^{k-1},X_{i}^{k+1},\dots,X_{i}^{N-1},\mathbf{M^{\prime}},G_{i})$
$\ \ \ \ +\frac{1}{2}\sum_{k,i}I_{{\bar{\mathbb{P}}_{\mathcal{A}\mathcal{B}\mathcal{E}}}}(X_{i}^{k}:X_{i}^{1},\dots,X_{i}^{k-1},X_{i}^{k+1},\dots,X_{i}^{N-1},\mathbf{M^{\prime}},G_{i})$
$\leq\sum_{k,i}I_{{{\mathbb{P}}_{\mathcal{A}\mathcal{B}\mathcal{E}}}}(X_{i}^{k}:X_{i}^{1},\dots,X_{i}^{k-1},X_{i}^{k+1},\dots,X_{i}^{N-1},\mathbf{M^{\prime}},G_{i}),$ (42)
where for the first inequality we have invoked the fact that the mutual information is a convex function of the probability distributions over the protocol-specific variables $\{X_{i}^{k}\},\{G_{i}\},\mathbf{M^{\prime}}$ which in turn are convex linear functions of the underlying tripartite no-signaling distributions ${\mathbb{P}}^{(\alpha,\gamma)},\bar{\mathbb{P}}^{(\alpha,\gamma)},\tilde{\mathbb{P}}^{(\alpha,\gamma)}$, while the second inequality follows from (36). Yet again, the terms on the right-hand side of (7) and (8) remain unaltered, implying that the distribution ~$\tilde{\mathbb{P}}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}$ of the desired form (9), satisfies the bipartite (7) and multipartite (8) formulation of $\mathcal{IC}$ while retaining the value of the CHSH expressions (27) (28). This completes the proof, and establishes the optimality of the slice (9).

### v2

@@ Abstract
$\beta(\mathcal{A},\mathcal{B})_{\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\text{iso}^{\prime}(\alpha,\gamma)}}=\frac{1+\alpha}{2};$ (28)
$\beta(\mathcal{B},\mathcal{E})_{\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\text{iso}^{\prime}(\alpha,\gamma)}}=\frac{1+\gamma}{2}.$ (29)
We are now prepared to address the central problem: Finding the maximum value of $\beta(\mathcal{B},\mathcal{E})$ given a fixed value of $\beta(\mathcal{A},\mathcal{B})$, under the constraints imposed by the information-theoretic criteria (6), (7), and (11) associated with the $\mathcal{IC}$ principle. This can be framed as the following optimization problem:
max $\quad\beta(\mathcal{B},\mathcal{E})_{\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}}$
subj to $\quad\mathcal{I}_{\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}}^{q}\leq\mathcal{C}^{q}$ (30)
$\quad\beta(\mathcal{A},\mathcal{B})=p$
for a fixed value $p\in[0.5,1]$. The $\mathcal{IC}$ constraints in Eqs. (6), (7), (8), and (11) are generically expressed as $\mathcal{I}_{\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}}^{q}\leq\mathcal{C}^{q}$, where $\mathcal{I}_{\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}}^{q}$ is the $\mathcal{IC}$ functional—i.e., a sum of mutual information terms depending on $\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}$—and $\mathcal{C}^{q}$ denotes constant parameters fixed by the protocol and communication channel. It is evident the $\mathcal{IC}$ functional $\mathcal{I}$ depends on the joint probability distribution, $P$, of all variables involved in the scenario, however, since the protocol is considered fixed for (Proof of Lemma 1), we may write $\mathcal{I}_{P}^{q}=\mathcal{I}_{\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}}^{q}$.
Suppose there is an optimal tripartite distribution, $\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{*}$, solving the optimization problem (Proof of Lemma 1). As we demonstrated, this distribution can be transformed via LOSR into $\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\text{iso}^{\prime}(\alpha,\gamma)}$, keeping the CHSH values, $\beta(\mathcal{A},\mathcal{B})$ and $\beta(\mathcal{B},\mathcal{E})$. In parallel, as LOSR constitutes a local post-processing operation, which, according to the data processing inequality (13) yields:
$\mathcal{I}_{\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\text{iso}^{\prime}(\alpha,\gamma)}}^{q}\leq\mathcal{I}_{\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{*}}^{q}.$ (31)
[…]
$\beta(\mathcal{B},\mathcal{E})_{\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{*}}\in\left\{\beta(\mathcal{B},\mathcal{E})_{\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\text{iso}^{\prime}(\alpha,\gamma)}}\right\}_{\alpha,\gamma}.$ (32)
We now turn to identifying the specific distributions $L_{\mathcal{A}}^{\prime*},L_{\mathcal{E}}^{*},\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\prime\prime\prime*}$ that solve the optimization problem in Eq.(Proof of Lemma 1). Let us first focus in $\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\prime\prime\prime*}$. In this case, we note that maximally uncorrelated distribution $W_{\mathcal{A}\mathcal{B}\mathcal{E}}$ ($p_{W_{\mathcal{A}\mathcal{B}\mathcal{E}}}(a,b,e|x,y,z)=1/8$) provides the minimum value for the $\mathcal{IC}$ functional, i.e., $\mathcal{I}_{W_{\mathcal{A}\mathcal{B}\mathcal{E}}}^{q}\leq\mathcal{I}_{\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}}^{q}$, $\forall\>\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}$. As a consequence, any other distribution, $\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\prime\prime\prime}$, different of $W_{\mathcal{A}\mathcal{B}\mathcal{E}}$ cannot yield on smaller values for the $\mathcal{IC}$ functional than the case $\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\prime\prime\prime}=W_{\mathcal{A}\mathcal{B}\mathcal{E}}$. This can be easily seen by invoking the convexity of mutual information on probability distributions, which considering (26) gives,
$\mathcal{I}_{\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\text{iso}^{\prime}(\alpha,\gamma)}}^{q}\;\;\leq\alpha\mathcal{I}_{\text{PR}_{\mathcal{A}\mathcal{B}}\otimes L_{\mathcal{E}}}^{q}+\gamma\mathcal{I}_{\text{PR}_{\mathcal{B}\mathcal{E}}\otimes L_{\mathcal{A}}^{\prime}}^{q}+(1-\alpha-\gamma)\mathcal{I}_{\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\prime\prime\prime}}^{q},$ (33a)
$\mathcal{I}_{\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\text{iso}^{\prime}(\alpha,\gamma)W}}^{q}\leq\alpha\mathcal{I}_{\text{PR}_{\mathcal{A}\mathcal{B}}\otimes L_{\mathcal{E}}}^{q}+\gamma\mathcal{I}_{\text{PR}_{\mathcal{B}\mathcal{E}}\otimes L_{\mathcal{A}}^{\prime}}^{q}+(1-\alpha-\gamma)\mathcal{I}_{W_{\mathcal{A}\mathcal{B}\mathcal{E}}}^{q},$ (33b)
where $\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\text{iso}^{\prime}(\alpha,\gamma)W}$ denotes the particular case of $\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\text{iso}^{\prime}(\alpha,\gamma)}$ when $\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\prime\prime\prime}=W_{\mathcal{A}\mathcal{B}\mathcal{E}}$ in (26). Subtracting both relations in (33) and considering $\mathcal{I}_{W_{\mathcal{A}\mathcal{B}\mathcal{E}}}^{q}\leq\mathcal{I}_{\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\prime\prime\prime}}^{q}$, we have
$\mathcal{I}_{\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\text{iso}^{\prime}(\alpha,\gamma)W}}^{q}\leq\mathcal{I}_{\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\text{iso}^{\prime}(\alpha,\gamma)}}^{q}.$ (34)
In parallel, from (27), it follows that the distribution $\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\prime\prime\prime}$ does not contribute to the CHSH values; that is, $\beta(\mathcal{A},\mathcal{B})_{\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\prime\prime\prime}}=\beta(\mathcal{B},\mathcal{E})_{\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\prime\prime\prime}}=1/2$. Combining this fact with (34), we conclude that any distribution $\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\prime\prime\prime}$, different from $W_{\mathcal{A}\mathcal{B}\mathcal{E}}$, performs at most as good as the case $\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\prime\prime\prime}=W_{\mathcal{A}\mathcal{B}\mathcal{E}}$. This is because (34) ensures that, in general, any such $\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\prime\prime\prime}$ brings the value of the $\mathcal{IC}$ functional closer to the upper bound $\mathcal{C}^{q}$—regardless of $\alpha$ and $\gamma$—without contributing for the increasing of $\beta(\mathcal{B},\mathcal{E})$. Consequently, any value $\beta(\mathcal{B},\mathcal{E})_{\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\text{iso}^{\prime}(\alpha,\gamma)}}$, such that $\beta(\mathcal{B},\mathcal{E})_{\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\text{iso}^{\prime}(\alpha,\gamma)}}>\max\{\beta(\mathcal{B},\mathcal{E})_{\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\text{iso}^{\prime}(\alpha,\gamma)W}}\}$, will result in a violation of $\mathcal{IC}$, as ensured by (34). Therefore, it is sufficient to take the optimal distribution $\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{*}$ as the following form:
$\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{*(\alpha,\gamma)}=\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\text{iso}^{\prime}(\alpha,\gamma)W}=\alpha\text{PR}_{\mathcal{A}\mathcal{B}}\otimes L_{\mathcal{E}}^{*}+\gamma\text{PR}_{\mathcal{B}\mathcal{E}}\otimes L_{\mathcal{A}}^{\prime*}+(1-\alpha-\gamma)W_{\mathcal{A}\mathcal{B}\mathcal{E}},$ (35)
which is a family of two-parameter slices of tripartite no-signaling polytope characterized by the local distributions $L_{\mathcal{A}}^{\prime*},L_{\mathcal{E}}^{*}$.
Next, we show that in order to derive monogamy relations of the form (2)—both for the bipartite cases (6), (7) and the multipartite criteria (8), (11) associated with $\mathcal{IC}$—the optimal choice for the local distributions is white noise: specifically, $L_{\mathcal{A}}^{\prime*}=W_{\mathcal{A}}$ and $L_{\mathcal{E}}^{*}=W_{\mathcal{E}}$, where $p_{\text{W}_{\mathcal{A}}}(a|x)=p_{\text{W}_{\mathcal{E}}}(e|z)=1/2$. To establish this claim, consider the distribution $\bar{\mathbb{P}}^{*(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}$, obtained from $\mathbb{P}^{*(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}$ by flipping all outputs. This leads to:
$p_{\bar{\mathbb{P}}^{*(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}}(a,b,e|x,y,z)=\alpha p_{\text{PR}_{\mathcal{A}\mathcal{B}}\otimes\bar{L}_{\mathcal{E}}^{*}}(a,b,e|x,y,z)+\gamma p_{\text{PR}_{\mathcal{B}\mathcal{E}}\otimes\bar{L}_{\mathcal{A}}^{*}}($ $a,b,e|x,y,z)$ (36)
$+(1-\alpha-\gamma)p_{W_{\mathcal{A}\mathcal{B}\mathcal{E}}}(a,b,e|x,y,z),$
where, $p_{\bar{L}_{\mathcal{E}}^{*}}(e|z)=p_{{L}_{\mathcal{E}}^{*}}(e\oplus 1|z)$ and $p_{\bar{L}_{\mathcal{A}}^{*}}(a|x)=p_{{L}_{\mathcal{A}}^{*}}(a\oplus 1|x)$. As flipping the outputs is a simple local post-processing, similarly to (31), data processing inequality (13) ensures that the $\mathcal{IC}$ functionals on bipartite (6),(7) and multipartite criteria (8),(11) should only decrease, i.e.:
$\mathcal{I}_{\bar{\mathbb{P}}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{*(\alpha,\gamma)}}^{q}\leq\mathcal{I}_{\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{*(\alpha,\gamma)}}^{q}.$ (37)
Moreover, as previously mentioned for the LOSR procedures, the upper bound, $\mathcal{C}^{q}$, remains unaltered. Hence, given a distribution, ${\mathbb{P}}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{*(\alpha,\gamma)}$, which satisfies a $\mathcal{IC}$ criterion of interest, $\mathcal{I}_{\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{*(\alpha,\gamma)}}^{q}\leq\mathcal{C}^{q}$, then the flipped-outcomes distribution, $\bar{\mathbb{P}}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{*(\alpha,\gamma)}$, also satisfies the same criterion $q$. Note, that the value of the CHSH expression, $\beta(\mathcal{A},\mathcal{B})$ and $\beta(\mathcal{B},\mathcal{E})$ remain unaltered when all outcomes are simultaneously flipped, as they only depend on the coefficients $(\alpha,\gamma)$ (28) (29).
Now, let us consider another tripartite distribution, $\tilde{\mathbb{P}}^{*(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}$, obtained by mixing equal proportions of the original tripartite distribution ${\mathbb{P}}^{*(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}$ and the flipped-outcomes distribution $\bar{\mathbb{P}}^{*(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}$, i.e., $\tilde{\mathbb{P}}^{*(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}=\frac{1}{2}({\mathbb{P}}^{*(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}+\bar{\mathbb{P}}^{*(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}})$, such that,
$p_{\tilde{\mathbb{P}}^{*(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}}(a,b,e|x,y,z)=\frac{1}{2}p_{\mathbb{P}^{*(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}}(a,b,e|x,y,z)+\frac{1}{2}p_{\bar{\mathbb{P}}^{*(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}}(a,b,e|x,y,z),$ (38)
@@
$p_{\tilde{\mathbb{P}}^{*(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}}(a,b,e|x,y,z)$ $=\alpha p_{\text{PR}_{\mathcal{A}\mathcal{B}}\otimes W_{\mathcal{E}}}(a,b,e|x,y,z)+\gamma p_{\text{PR}_{\mathcal{B}\mathcal{E}}\otimes W_{\mathcal{A}}}(a,b,e|x,y,z)$ (39)
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+(1-\alpha-\gamma)p_{W_{\mathcal{A}\mathcal{B}\mathcal{E}}}(a,b,e|x,y,z),$
where we have used the facts, $\frac{1}{2}(L_{\mathcal{B}}^{*}+\bar{L}_{\mathcal{B}}^{*})=W_{\mathcal{B}}$, $\frac{1}{2}(L_{\mathcal{E}}^{*}+\bar{L}_{\mathcal{E}}^{*})=W_{\mathcal{E}}$, $\frac{1}{2}\left(W_{\mathcal{A}\mathcal{B}\mathcal{E}}+\bar{W}_{\mathcal{A}\mathcal{B}\mathcal{E}}\right)=W_{\mathcal{A}\mathcal{B}\mathcal{E}}$. Notice that the values of the CHSH expression, $\beta(\mathcal{A},\mathcal{B})$ and $\beta(\mathcal{B},\mathcal{E})$ still remain unaltered as they only depend on the coefficients $(\alpha,\gamma)$ in (28) and (29). At this point, we invoke convexity of mutual information once more, which allows to write for the $\mathcal{IC}$ functional considering (38) :
$\mathcal{I}_{\tilde{\mathbb{P}}^{*(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}}^{q}\leq\frac{1}{2}\mathcal{I}_{\mathbb{P}^{*(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}}^{q}+\frac{1}{2}\mathcal{I}_{\bar{\mathbb{P}}^{*(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}}^{q}.$ (40)
Thus, if the distribution ${\mathbb{P}}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{*(\alpha,\gamma)}$ satisfies a $\mathcal{IC}$ criterion of interest, $\mathcal{I}_{\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{*(\alpha,\gamma)}}^{q}\leq\mathcal{C}^{q}$, the distribution, $\tilde{\mathbb{P}}^{*(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}$, obtained from (38) will necessarily satify it, having the same value of the CHSH expressions (28) and (29) as ${\mathbb{P}}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{*(\alpha,\gamma)}$. Consequently, any optimal distribution, ${\mathbb{P}}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{*(\alpha,\gamma)}$, solving the problem (Proof of Lemma 1), characterized by the optimal local distributions $L_{\mathcal{A}}^{\prime*}$ and $L_{\mathcal{E}}^{*}$, will produce the same CHSH values as $\tilde{\mathbb{P}}^{*(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}$, while $\tilde{\mathbb{P}}^{*(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}$ also satisfies $\mathcal{IC}$. Thus, in order to find the optimal CHSH values as permitted by $\mathcal{IC}$ criteria in the problem (Proof of Lemma 1) it is enough to analyze the slice (39), which assumes the form presented in the main text:
@@
$p_{{\mathbb{P}}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{*(\alpha,\gamma)}}(a,b,e|x,y,z)=p_{\tilde{\mathbb{P}}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}}(a,b,e|x,y,z)=\alpha\frac{1}{4}\delta_{a\oplus b,xy}+\gamma\frac{1}{4}\delta_{b\oplus e,yz}+(1-\alpha-\gamma)1/8.$ (41)
This completes the proof, and establishes the optimality of the slice (9).